OpenStudy (anonymous):

y'''-2y''-2y'+y=0 where y"(0)=1, y(0)=2, y(0)=2 how to get this were having trouble in this assignment please help :(

OpenStudy (anonymous):

First, solve the characteristic equation \[ r^3-2 r^2-2 r+1=0\\ (1 + r) (1 - 3 r + r^2)=0 \] The roots are \[ \{r= -1\},\\ \left\{r= \frac{1}{2} \left(3-\sqrt{5}\right)\right\}, \\\left\{r= \frac{1}{2} \left(3+\sqrt{5}\right)\right\} \]

OpenStudy (anonymous):

You find the general solution \[ y=a e^{\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) x}+b e^{\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) x}+c e^{-x} \]

OpenStudy (anonymous):

You find a, b and c from the initial conditions.

OpenStudy (anonymous):

thank you @eliassaab

OpenStudy (anonymous):

i'm now confuse how to find those value :(

OpenStudy (anonymous):

i need clear solution please im begging for help

OpenStudy (anonymous):

sorry if im too slow to understand...

OpenStudy (anonymous):

You have three initial conditions and you need three equations, so from each condition get an equation. From y(0)=2, you get \[ 2=y(0)=a e^{\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) 0}+b e^{\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) 0}+c e^{-0}=a+b+c=2 \] This is the first equation between a, b and c

OpenStudy (anonymous):

\[ y'(x)=\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) a e^{\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) x}+\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) b e^{\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) x}-c e^{-x}\\ y'(0)=\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) a+\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) b-c=2 \] This is another equation between a, b and c

OpenStudy (anonymous):

\[ y''(x)=\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right)^2 a e^{\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) x}+\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right)^2 b e^{\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) x}+c e^{-x}\\ y''(0)=\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right)^2 a+\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right)^2 b+c=1 \] This is the third equation. Is your teacher a masochist? I refuse to solve these three eqautins by hand for a , b and c

OpenStudy (anonymous):

sometimes but shes kind i guest???

OpenStudy (anonymous):

I solved them using Mathematica and I got \[ a= \frac{1}{10} \left(13+5 \sqrt{5}\right)\\b= \frac{13 \sqrt{5}-25}{10 \sqrt{5}}\\c= -\frac{3}{5} \]

OpenStudy (anonymous):

Finally the answer is \[ y=\frac{1}{10} \left(13+5 \sqrt{5}\right) e^{\left(\frac{3}{2}-\frac{\sqrt{5}}{2}\right) x}+\frac{\left(13 \sqrt{5}-25\right) e^{\left(\frac{3}{2}+\frac{\sqrt{5}}{2}\right) x}}{10 \sqrt{5}}-\frac{3 e^{-x}}{5} \]

OpenStudy (anonymous):

I would never assign such a problem to be done by hand.

OpenStudy (anonymous):

I have to go. It was nice helping you.

OpenStudy (anonymous):

thank you soo much till next time

OpenStudy (anonymous):

YW